### v1

# SimuHome: A Temporal- and Environment-Aware Benchmark for Smart Home LLM Agents

Abstract
Large Language Model (LLM) agents excel at multi-step, tool-augmented tasks. However, smart homes introduce distinct challenges, requiring agents to handle latent user intents, temporal dependencies, device constraints, scheduling, and more. The main bottlenecks for developing smart home agents with such capabilities include the lack of a realistic simulation environment where agents can interact with devices and observe the results, as well as a challenging benchmark to evaluate them. To address this, we introduce SimuHome, a time-accelerated home environment that simulates smart devices, supports API calls, and reflects changes in environmental variables. By building the simulator on the Matter protocol [^1] https://csa-iot.org/all-solutions/matter/—the global industry standard for smart home communication—SimuHome provides a high-fidelity environment, and agents validated in SimuHome can be deployed on real Matter-compliant devices with minimal adaptation. We provide a challenging benchmark of 600 episodes across twelve user query types that require the aforementioned capabilities. Our evaluation of 11 agents under a unified ReAct framework reveals that while models perform well on simple tasks, they struggle with latent intent inference, state verification, and especially temporal scheduling. Even the top-performing model, GPT-4.1, reaches only 54% success rate. These findings highlight a critical need for methods that can reliably verify the current state via tools before acting and coordinate time-dependent actions. [^2] We will release our code and dataset upon publication of the paper.
1 Introduction
Recently, Large Language Model (LLM) agents have demonstrated strong abilities on multi-step, tool-augmented tasks, including API retrieval, invocation, and intermediate state verification (Qin et al., 2024; Patil et al., 2025; Chen et al., 2024; Huang et al., 2024; Xu et al., 2023; Schick et al., 2023). These abilities enable long-horizon tasks such as web navigation and goal pursuit, where agents must plan, check states, and validate outcomes over multiple steps (Zhou et al., 2024; Yao et al., 2022; Deng et al., 2023; Xie et al., 2024; Yao et al., 2024; Trivedi et al., 2024).
Smart home agents, such as Amazon Alexa and Google Home, are among the earliest productionized tool agents in the real world and have long been a research topic. To meet real-world challenges, smart home agents need capabilities to handle many factors, such as: (1) latent user intents (e.g., “It feels stuffy” implying humidity control), (2) temporal dependencies (e.g., “Turn on the kitchen light when the dishwasher finishes”), (3) dependencies among device actions and attributes (e.g., a dishwasher cannot be opened while it is running), (4) scheduling (e.g., “Play music in the morning”). However, most (if not all) smart home agents to date fall short in all these areas. One of the critical bottlenecks is the lack of training and test data with such complexities. Even if such datasets existed, static datasets have clear limitations: agents cannot learn by doing, and agent performance cannot be evaluated accurately (because a user intent may be satisfied in multiple ways that are not annotated in the dataset). We aim to address this challenge by developing a high-fidelity smart home simulator in which agents can interact with devices through APIs and observe the results reflected in the environment, along with an extensive benchmark containing a variety of complex user requests, both feasible and infeasible.
Figure 1: The SimuHome home environment with Matter-compliant devices, featuring a GUI where users can arrange devices across rooms, configure their attributes, and evaluate agent reasoning for multi-device control.
Our first contribution is a smart home simulator, SimuHome (Figure 1). SimuHome is a time-accelerated smart home environment that accommodates various room layouts, environmental variables (e.g., temperature, illuminance), and smart devices. Agents can call APIs to operate devices (e.g., set the AC to 25 degrees). Devices are simulated with internal constraints checked (e.g., the AC must be turned on to set its temperature), and the results affect the environment (e.g., the room temperature gradually drops to 25 degrees over 10 minutes). Notably, SimuHome implements Matter, a broadly adopted smart-home interoperability standard. As a result, the attributes and constraints within devices are high fidelity. Moreover, agents trained and verified in SimuHome can run on real Matter-compliant devices with minimal adaptation. SimuHome also enables controlled experiments in a cheap and fast way. It allows unlimited experimentation, including stress-testing rare edge cases and counterfactual scenarios, while strict reproducibility ensures fair comparisons and iterative validation across models. Although beyond the scope of our work, it can also support model training through reinforcement learning.
Our second contribution is a manually validated benchmark of 600 episodes covering twelve user query types, each provided in feasible and infeasible variants to assess agents’ abilities in proactive intent inference, dynamic state and physical-limit checks, and temporal scheduling. Each case is packaged as a single episode with an initial home state (i.e., rooms, device states, environmental variables), a verifiable goal, a natural-language query, and a set of required actions that enforce information gathering before control. Feasible cases are scored by comparing the resulting state in SimuHome with the target state. Infeasible cases, which embed false premises, physical limits, or temporal conflicts, are assessed by LLM judges.
We evaluate 11 LLM agents under a unified ReAct (Yao et al., 2023) setup across 600 episodes with feasible and infeasible variants, scoring feasible tasks by simulator state comparisons and assessing infeasible logic checks with validated LLM judges. Models handle simple retrieval and explicit device control comparatively well. They often fail to infer latent intent or to verify current device and environment states before acting. Temporal scheduling is the most challenging area: contradiction blindness and mishandling are common, and even GPT-4.1 reaches only 54% success rate. These results motivate methods that verify the current state via tools before acting and that plan and coordinate time-dependent actions reliably in dynamic home environments.
2 Related Work
Simulated Benchmarks for Household Embodied Agents.
Embodied-agent benchmarks have advanced instruction following in household settings, but interactions with devices are usually limited to oversimplified actions that overlook real-world constraints. AI2-THOR (Kolve et al., 2017) enables agents to navigate photorealistic 3D rooms and manipulate objects through atomic actions (e.g., open/close, pick up/put down). ALFRED (Shridhar et al., 2020) extends this to long-horizon tasks, requiring agents to translate language and first-person observations into action sequences that yield persistent state changes, supported by $\sim$25k demonstrations. VirtualHome (Puig et al., 2018) captures everyday activities (e.g., cooking dinner, cleaning a room) as executable programs derived from crowdsourced scripts. While effective for language grounding and task structure, these simulators constrain devices to discrete commands (ToggleOn/Off, Open/Close), missing communication delays, conflicts, and cascading cross-device effects that arise in real homes.
LLM Agents and Benchmarks for Smart Homes.
Recent smart home LLM benchmarks emphasize planning and goal interpretation but similarly rely on simplified abstractions. HomeBench (Li et al., 2025) evaluates instruction following under valid, invalid, and mixed requests across single- and multi-device settings, highlighting error detection, refusal, and coordinated execution. Sasha (King et al., 2024) studies goal interpretation, mapping underspecified intentions to device-level plans and assessing their quality via user studies. SAGE (Rivkin et al., 2023) frames smart home control as sequential tool use, guiding LLMs through API calls, preference handling, and state monitoring. Despite these advances, current suites operate in pre-scripted environments and omit dynamic device attributes or temporal constraints, limiting their fidelity to real households.
SimuHome addresses this gap with a reproducible simulator that models device effects on ambient conditions while supporting attribute tracking, precondition enforcement, and temporal constraint handling.
3 SimuHome: A Smart Home Simulator
3.1 Motivation
Evaluating LLM agents in a smart home requires a simulator that mirrors the real world’s continuous and reactive nature. However, existing simulators for agents have a limitation. They do not simulate the realistic chain reaction where one action can affect others and the environment; instead, each command is treated as a separate, isolated event. To address this problem, we design SimuHome around four core requirements:
Complex Temporal Constraints. To evaluate an agent’s temporal reasoning, the simulator must handle a variety of complex time-based queries (e.g.,“Keep the kitchen lights on until the dishwasher finishes”). This allows us to test if the agent can understand and plan actions with complex temporal dependencies.
Dependency Modeling Based on an Industry Standard. The simulator realistically models the operational rules of smart devices according to the Matter industry standard. This design allows us to evaluate whether the agent can learn and adapt to real-world device constraints. For example, the simulator enforces the rule that an air conditioner’s power must be on before its fan speed can be changed, enabling us to test if the agent understands this dependency.
Real-Time Environmental Feedback. The simulator models the continuous, real-time effects of device actions on the environment (e.g., temperature and illuminance). This creates a dynamic setting to test if the agent can monitor ongoing changes and react appropriately, rather than just acting on static information. For example, as an air conditioner runs, the temperature gradually drops, and the agent must perceive this change to complete its goal.
Reproducibility. The environment must be perfectly reproducible, ensuring that an agent’s actions produce identical outcomes under the same initial conditions. This is crucial for reliably measuring and comparing the performance of different agents or strategies.
3.2 Simulator Architecture and Operation
Our simulator operates by processing time in fixed intervals. The fundamental unit of time, a tick, is defined as 0.1 real-world seconds. All environmental and device state updates are calculated at every tick. This method of updating the state at a fixed interval allows the simulator to model the outcomes of processes that occur continuously in the real world with high fidelity. The simulator comprises three components: the Smart Home Environment, the Real-Time State Update Mechanism, and the Agent-Simulator Interface.
Smart Home Environment. A home is a configurable environment composed of one or more rooms, each containing a custom set of devices and four environmental variables: temperature, illuminance, humidity, and air quality. To enable realistic scenarios, the environment includes both devices that directly influence environmental variables (e.g., an air conditioner) and those with multi-stage operational cycles (e.g., a washing machine). In total, we model 17 distinct device types. A full list of these devices can be found in Appendix A.4.
Real-Time State Update Mechanism. The core of the simulation is the Aggregator module, which models the dynamic impact of device operations on the environment. At each tick, the Aggregator calculates the combined influence of all active devices on their relevant environmental factors. For example, temperature is affected by air conditioners and heat pumps, illuminance by lights, humidity by humidifiers/dehumidifiers, and air quality by air purifiers. The magnitude of this influence is cumulative; it scales with the number of active devices and their specific settings (e.g., the fan speed of an air conditioner). This mechanism ensures that the environment responds realistically to agent actions. The detailed update equations for the Aggregator are provided in Appendix A.12.
Agent-Simulator Interface. The agent interacts with the simulator by invoking a set of 13 tools. The structure of these tools mirrors Matter’s modular approach to defining device capabilities. Detailed tool specifications are provided in Appendix A.2.
3.3 Task Definition
SimuHome tasks are modeled as a partially observable Markov decision process (POMDP) $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},\mathcal{R})$. The environment state $s_{t}\in\mathcal{S}$ consists of the device state, represented by the Matter hierarchical model of Endpoints, Clusters, and Attributes, and the environmental state, defined by ambient conditions such as temperature, illuminance, humidity, and air quality. At each tick, the agent executes an action $a_{t}\in\mathcal{A}$, implemented as a Matter Command, which updates the device state. The transition function $\mathcal{T}$ applies the Aggregator mechanism to propagate device effects onto the environmental state. The agent receives an observation $o_{t}\in\mathcal{O}$, corresponding to the subset of device attributes and environmental state variables exposed through the API, which provides only partial visibility into the full state. The reward function $\mathcal{R}$ is defined as part of the evaluation process given a task query. Details of how rewards are assigned are provided in §5.1.
4 Benchmark Design
4.1 Query Types
We define twelve query types that commonly arise in user queries within smart home environments. These are designed to evaluate an agent’s abilities in device control, environmental variable queries such as temperature and illuminance, implicit intent inference, and temporal coordination with three sub-types. Each type is paired with an infeasible scenario to test the agent’s capacity for logical consistency and constraint handling, yielding a total of 12 categories. See Appendix A.1 for examples of infeasible scenarios corresponding to each query type.
QT1 (Environment Perception). This evaluates the ability to correctly perceive environmental conditions and device statuses, and then provide accurate, logical information in natural language. For example, in response to “I’m about to cook, can you tell me how humid it is in the kitchen?”, the agent must identify the kitchen area, use an environment-query tool to check the humidity, and respond with clear units and values. If device discovery is needed during this process, the agent must first check the list of devices in that room.
QT2 (Implicit Intent). This assesses the ability to infer the user’s underlying goal from complaints or indirect expressions and to create and execute a suitable device control plan to address it. For instance, upon hearing “It feels too stuffy here in the living room”, the agent should check the living room’s humidity and then take action to adjust it, such as turning on a humidifier or turning off a dehumidifier.
QT3 (Explicit Intent). This evaluates the ability to accurately interpret and execute commands involving specified devices and target values. For example, for the command “Set the living room air purifier fan speed to one hundred percent, the strongest power”, the agent must verify the presence of an air purifier in the living room. If it is off, the agent must turn it on first before setting the fan speed to 100%.
QT4-1 (Future Scheduling). This assesses the ability to schedule and plan the control of multiple devices (e.g., lights, air conditioners) to activate at a specific future time. For example, for the request “I will go to sleep in ten minutes. Can you turn off the lights and the humidifier in ten minutes?”, the agent must calculate the absolute time 10 minutes from the current time. It should then schedule both actions as a single, conflict-free workflow. Before registering the commands, the agent must pre-validate that each device is controllable and the specified parameters are within acceptable ranges.
QT4-2 (Dependency Scheduling). This evaluates the ability to create a coordinated schedule for an operational device (one that takes time to complete, such as a dishwasher) and an instantaneous device such as a light, considering dependencies and completion times. For example, for the request “When the dishwasher finishes, please turn off the kitchen lights”, the agent must check the dishwasher’s remaining operating time to calculate its estimated completion time. It should then schedule the lights to turn off based on that absolute time, after verifying and registering the correct parameters and sequence for the command.
QT4-3 (Concurrent Scheduling). This assesses the ability to schedule two or more operational devices to work without conflict, according to given time constraints. For example, for the request “Schedule the dishwasher so that it completes at the same time the washer finishes”, the agent must check the remaining operating time of both devices to calculate if a simultaneous finish is possible. If it is, the agent should adjust the start time of one device and register a workflow to ensure they finish together.
4.2 Episode Generation
Figure 2: Episode Generation Pipeline
Figure 3: Episode Evaluation Pipeline
Definition and Components of Episode. An episode defines a single, self-contained task scenario for the agent. As illustrated in Figure 2, each episode is composed of four key components: the initial home state (including room layouts, device states, and environmental variable values), a goal the agent must achieve, the natural language user query, and the set of required actions for evaluation.
STEP1: Initial Home State Construction. The initial home state for each episode is constructed in two stages to ensure diverse and realistic starting conditions (Figure 2). First, a variety of physical layouts with different room and device configurations are generated. Second, starting from an all-off state, devices are operated randomly, establishing plausible device states. Although this process involves randomization, it is controlled by a seed to ensure that both the layout and the initial state are fully reproducible.
STEP2: Goals and Required Actions. A goal defines the desired final state of specific devices or environmental variables that the agent must achieve. The generation process, which varies by query type (see Appendix A.5), is designed to ensure all goals are logically consistent. For instance, as illustrated for QT3 in Figure 2 (Step 3), a device goal is created by sampling from a pre-defined set of valid states (e.g., onoff: on, fanspeed: 40%). Each of these state sets is constructed to inherently satisfy the device’s internal dependencies. Required Actions are a sequence of tool calls that an agent must perform. This ensures the agent’s subsequent actions are based on up-to-date information gathered from the environment. For example, before attempting to change an air purifier’s fan speed, the agent is required to first invoke the tool get_room_devices(utility_room) to confirm the device’s existence. An episode is marked as successful only if the agent both satisfies the goal and its tool call history contains all required actions.
STEP3: Query Synthesis. In general, a user’s natural language query embodies a goal to be achieved, and the clarity of this goal is essential for an accurate evaluation of the agent’s success. Therefore, we first defined a verifiable goal for the agent to accomplish and subsequently synthesized a natural language query based on it. We then used GPT-5-mini (OpenAI, 2025b) to synthesize the natural language queries from these predefined goals. To ensure each query accurately reflected its predefined goal, two graduate students researching tool agents independently reviewed the entire dataset. Their inter-annotator agreement, measured using Cohen’s $\kappa$ coefficient (Cohen, 1960), was 0.92 for identifying queries that required correction. This demonstrates that the validation procedure for our dataset is highly consistent and reliable, suggesting that the benchmark data is composed of high-quality natural language queries.
STEP4: Episode Generation. By integrating the components generated in the preceding steps, we constructed our final benchmark dataset. We generated 50 distinct episodes for each of the 12 query types, resulting in a high-quality dataset of 600 episodes designed for evaluating smart home agents.
5 Evaluation
5.1 Evaluation Methods
As illustrated in Figure 3, we evaluate agent performance across the 12 query types defined in §4.1 using two complementary methods: simulator-based and LLM-judge-based evaluation.
Simulator-based Evaluation. Simulator-based evaluation is essential for episodes that target physical state changes because outcomes must be assessed objectively and reliably. At the end of each episode, the simulator automatically verifies the final states of all relevant devices and environmental variables, and compares them with the goal defined for that episode. In the QT4-Feasible episode shown in Figure 3, after the agent completes all its actions, the simulator accelerates time to 19:00, when the laundry cycle finishes. Subsequently, the goal state—the washer being off and the kitchen light being on—is compared with the simulator’s final state to determine success. This direct state comparison yields a clear and fully automated success criterion and enables fair model-to-model comparisons under the same conditions.
LLM-judge-based Evaluation. We employ an LLM-based judge for episodes where success depends not on physical state changes, but on the agent’s final natural-language response. This provides an objective and effective complement to simulator-based evaluation. In our procedure, the judge is provided with the episode goal and the user query, along with a concise description of any infeasible conditions that must be verified. We also provide the agent’s full ReAct trajectory so that the judge can verify whether the final answer is supported by a coherent and sufficient chain of reasoning. For example, in the QT4-Infeasible episode shown in Figure 3, the LLM-judge assesses the agent’s explanation of a scheduling conflict. The judge scores only the final user-facing answer while using the trajectory to confirm consistency. For each case, we asked the judge three times and adopted the consistent outcome to ensure reliability (Taubenfeld et al., 2025). Detailed prompt templates and rubrics are available in Appendix A.9. To validate our LLM-judges, we confirmed that their assessments achieved substantial agreement (Cohen’s $\kappa$ = 0.826) with human evaluations, which themselves showed very high inter-rater reliability (see Appendix A.6 for the full analysis).
5.2 Experiments
Experimental Setup. We quantitatively compare models’ inference, planning, and tool-calling performance across the 12 query types defined in §4.1. The evaluation covers 11 models spanning open-source (Llama, Qwen, Gemma) and closed-source (Gemini, GPT) families, with reproducibility details provided in Appendix A.7. We employ the ReAct framework (Yao et al., 2023) for all experiments, enabling the models to generate both reasoning traces and actions in a step-by-step manner within our smart home simulation environment. The specific agent prompt used for the ReAct framework is provided in Appendix A.8.
5.3 Main Results
Table 1: Evaluation results show success rates (in %) across query types (QTs). F and IF refer to Feasible and Infeasible episodes, respectively. Superscripts J{J} and S{S} indicate results from LLM-judge-based and simulator-based evaluation, respectively.
QT1 QT2 QT3 QT4-1 QT4-2 QT4-3
Model FJ IFJ FS IFJ FS IFJ FS IFJ FS IFJ FS IFJ
Llama-4-Scout $58$ $42$ 22 $22$ $24$ $34$ 44 44 22 22 22 0
Llama-4-Maverick 96 $78$ $52$ $36$ $88$ $74$ $22$ $14$ $18$ $10$ 32 88
Qwen3-32B $82$ $66$ 62 $30$ $52$ $68$ $18$ $14$ $14$ 88 $16$ 66
Qwen3-235B-A22B $86$ $74$ $32$ $36$ 84 $70$ 26 $18$ $38$ $34$ $28$ $48$
Gemma-3-12B-it $78$ $38$ $14$ $32$ $32$ $24$ 22 0 0 0 0 0
Gemma-3-27B-it $80$ $48$ $54$ $24$ $48$ $44$ 44 22 $10$ 88 0 66
Gemini-2.5-Flash-Lite $78$ $60$ $44$ 50 $50$ $50$ 88 $34$ $10$ $16$ $16$ 2020
Gemini-2.5-Flash $92$ $86$ $66$ $54$ $82$ $74$ $22$ $44$ 40 32 $12$ 32
GPT-4.1-nano $58$ $42$ 66 $12$ $30$ $16$ 22 66 66 0 0 0
GPT-4.1-mini 96 $76$ 62 $28$ $64$ 76 26 40 40 2020 $10$ $28$
GPT-4.1 $98$ 82 $44$ $44$ 84 $88$ $50$ $12$ $46$ $34$ $34$ 32
Table 1 reports performance across all query types (QT1-QT4) under both feasible (F) and infeasible (IF) conditions. Several consistent patterns emerge.
Environment perception (QT1). QT1-F episodes are largely solved: Most frontier models exceeded 85%. In QT1-IF, we observe a success rate degradation across models, reflecting a consistent challenge in detecting requests based on false premises, such as references to non-existent devices.
Explicit vs. implicit device control (QT2 vs. QT3). Across both feasible and infeasible settings, most models handle explicit commands (QT3) substantially better than implicit requests (QT2). In feasible episodes, leading models exceed 84% on QT3, whereas QT2 plateaus at 62-66%. In feasible settings, models consistently perform better when users issue explicit commands (QT3) compared to implicit requests (QT2). In infeasible settings, however, the contrast becomes more pronounced: models can often detect nonexistent devices in QT3-IF with relatively high success rate, but they struggle far more when the device exists yet the goal is unattainable due to operational constraints. For example, an air conditioner already operating at maximum fan speed cannot lower the temperature further. This indicates that reasoning over such constraint-driven infeasibility poses a greater challenge for current models.
Temporal reasoning (QT4). QT4 queries, which require agents to handle scheduling, coordination, and temporal consistency, were the most challenging across both feasible and infeasible settings. GPT-4.1 achieved 50% on QT4-1, 46% on QT4-2, and 34% on QT4-3, outperforming peers but still far from robust. In infeasible cases with contradictory or impossible timing, performance degraded sharply: even GPT-4.1 and Gemini-2.5-Flash peaked at 44% (QT4-1-IF) and 34% (QT4-2-IF), with Qwen3-235B-A22B delivering notable performance at 48% on QT4-3-IF. These results highlight that contradiction detection and temporal validation remain open challenges.
Model-level comparison. Closed-source models (GPT-4.1, Gemini-2.5-Flash) consistently led overall, but none achieved stable performance across both feasible and infeasible tasks. Mid-sized open-source models such as Qwen3-235B-A22B and Gemini-2.5-Flash-Lite produced notable results on certain query types—for instance, Qwen3-235B-A22B reached 48% on QT4-3-IF—but their performance was uneven across query types. Smaller models (Gemma-3-12B-it, GPT-4.1-nano) remained limited overall, particularly on temporally demanding tasks.
Table 2: Error taxonomy. Detailed descriptions and examples are provided in Appendix A.10.
Category Error Type Definition
Feasible Environment Perception (EP) Failure to correctly perceive environmental variables.
Intent Inference (II) Misinterpreting the user’s underlying goal.
Device Control (DC) Operating the wrong device or command.
Action Planning (AP) Incomplete or incorrect planning of actions.
Temporal Reasoning (TR) Miscalculating times or sequence alignment.
Infeasible Contradiction Mishandling (CM) Detects a contradiction but fails to follow the instruction.
Contradiction Blindness (CB) Fails to detect a contradiction.
LLM-Judge (LJ) Misclassification by LLM-Judge.
Figure 4: Error type distributions of GPT-4.1 on QT2 and QT4. (a) and (b) show the distributions for feasible episodes, while (c) and (d) present the distributions for infeasible episodes.
6 Analysis
6.1 Error Analysis
We define eight error types to analyze agent failures: five for feasible episodes (EP, II, DC, AP, TR) and three for infeasible episodes (CM, CB, LJ). Error taxonomy is provided in Table 2. Our analysis centers on GPT-4.1, the best-performing model.
Figure 4 summarizes the error type distributions for GPT-4.1 across feasible and infeasible episodes. For feasible episodes, figure (a) and (b) show error distribution in QT2 and QT4. In QT2 (indirect requests), failures were dominated by Device Control (DC, 71%), where the model issued heuristic guesses instead of using the correct API. Intent Inference (II) errors (11%) also appeared, reflecting difficulty in mapping vague complaints such as “The room is too hot” to the appropriate device action. QT4 (temporal scheduling) exhibited a more diverse mix: DC (40%), Temporal Reasoning (TR, 25%), and Action Planning (AP, 19%) all contributed substantially, alongside smaller II errors (11%). These distributions show that multi-step temporal reasoning requires coordinating multiple skills simultaneously, making it substantially harder than direct execution tasks.
For infeasible queries, figure (c) and (d) highlight two dominant patterns. In QT1-QT3, GPT-4.1 often detected the contradiction but failed to follow the instructed protocol, resulting in Contradiction Mishandling (CM). For example, when asked to raise the kitchen temperature using a non-existent heat pump, it instead acted on the living-room heat pump. In QT4, the dominant issue was Contradiction Blindness (CB): the model failed to recognize temporal infeasibility (e.g., contradictory deadlines) and proceeded as if the request were valid. Even when contradictions were recognized, responses were frequently mishandled (CM).
6.2 Role of Tool Feedback
Figure 5: Tool-call error patterns of four models on QT3-F. The left chart shows the average number of errors relative to the average number of tool calls in successful cases. The right chart shows the proportion of tasks achieved through first-try success versus those requiring error recovery.
To better understand agent dynamics, we examined QT3, where most models were relatively strong. Figure 5 shows that over 40% of successful QT3 episodes involved recovery after an initial invalid tool call. In other words, agents did not require perfect prior knowledge of the Matter protocol but learned reactively from error messages. This ability to recover explains their robustness on explicit device-control queries. In contrast, the weakness on QT4 stems in part from its deferred-feedback: agents typically call the tool schedule_workflow, which returns only a scheduling acknowledgment (i.e., a success/failure message) without validating executability. Consequently, the simulator provides little corrective signal, leaving the agent unable to revise its plan.
7 Conclusion
We propose SimuHome, a Matter-aligned simulator and benchmark that reproducibly evaluates smart home LLM agents under realistic, dynamically changing conditions. We model 4 environmental variables (i.e., temperature, illuminance, humidity, air quality) and 17 device types with time-based effects and strict reproducibility, enabling near drop-in transfer to real Matter-compliant devices. We provide 600 episodes across 12 query types with feasible and infeasible variants, packaging each episode with an initial state, a verifiable goal, a natural-language query, and required actions for process-aware, objective scoring. We score feasible tasks by final state-to-goal comparison in the simulator and assess infeasible logic checks with LLM judge that shows high agreement with human evaluation. We evaluate 11 agents under the ReAct setup. Current open- and closed-source LLMs handle explicit control and simple retrieval well but often fail at latent-intent inference, live-state verification, and temporal scheduling; GPT-4.1 achieves an overall success rate of 54% across all query types.
Appendix A Appendix
A.1 Infeasible Query Types
QT1 Infeasible.
This evaluates the ability to identify requests based on a false premise, such as asking for information about non-existent devices or unsupported attributes. For example, for the request “Can you tell me the vendor ID for the air purifier in the living room?”, the agent must check the list of devices in the living room, confirm the absence of an air purifier, and explain that the request’s premise is invalid.
QT2 Infeasible.
This assesses the ability to identify situations where, even if the user’s intent is correctly inferred, the goal is impossible to achieve due to environmental constraints or device limitations. For example, in response to “The living room feels like a sauna”, the agent must verify that the living room’s cooling system is already operating at maximum capacity and explain, with supporting reasons, why further cooling is not possible.
QT3 Infeasible.
This evaluates the ability to identify and reject a command to control a non-existent device. For example, for the request “Turn on the humidifier in the living room”, the agent must check the device list for the living room and confirm the absence of a humidifier. It should then explain that the request cannot be fulfilled and terminate the task without altering any device’s state.
QT4-1 Infeasible.
This assesses the ability to identify and explain situations where a scheduling request is invalid because the user’s specified relative and absolute times are contradictory, or because the user has a misunderstanding of the current time. For example, if a user asks, “It’s 6 p.m. now, right? Turn on the kitchen light five minutes later at 6:05 p.m.”, but the actual time is not 6 p.m., the agent must check the current time, detect the discrepancy between the relative expression “five minutes later” and the absolute time “6:05 p.m.”, and clearly explain the contradiction.
QT4-2 Infeasible.
This evaluates the ability to identify and explain, with evidence, requests where the user incorrectly assumes a device’s completion time or creates a contradiction by providing both relative and absolute times. For example, suppose a washer is set to finish at 6:30 p.m., but the user requests, “I think the washer finishes at 6 p.m., so start the dehumidifier at 5:50 p.m., which is 10 minutes before it finishes”. The agent must check the washer’s actual estimated completion time. It then needs to point out that the user’s assumption (6 p.m.), the relative expression (“10 minutes before”), and the absolute time (“5:50 p.m.”) are all inconsistent. The agent must not register the schedule until the contradiction is resolved and should ask the user to reconfirm the correct timing.
QT4-3 Infeasible.
This evaluates the ability to identify and explain that a requested deadline is physically impossible to meet, given the current progress of two operating devices. For example, if the user requests, “Guests arrive at 6 p.m., so ensure both the washer and the dishwasher are completed by 5:30 p.m.”, the agent must check the current time and the minimum time required for each device to finish. Based on this, it should explain with clear reasoning why a 5:30 p.m. completion is not feasible and suggest the earliest possible completion time or an alternative sequential plan.
A.2 List of Tools
Table 3: Tool List for Agent
Name Description Args
finish Complete the task and return the final natural-language answer. answer (str, required): Final response text.
execute_command Execute a command on a device (e.g., turn on light, set level, set setpoint). device_id (str, req); endpoint_id (int, req); cluster_id (str, req); command_id (str, req); args (dict, req).
write_attribute Directly set a device attribute value. device_id (str, req); endpoint_id (int, req); cluster_id (str, req); attribute_id (str, req); value (any, req).
get_all_attributes Get all attributes of a device. device_id (str, req).
get_attribute Get a specific attribute of a device. device_id (str, req); endpoint_id (int, req); cluster_id (str, req); attribute_id (str, req).
get_device_structure Get device structure (endpoints, clusters, attributes, and commands). device_id (str, req).
get_rooms Get all rooms in the home along with their display names. (none)
get_room_devices Get all devices in a room. room_id (str, req).
get_room_states Get environmental states of a room (temperature, humidity, illuminance, PM10). room_id (str, req).
get_cluster_doc Perform semantic search across Matter cluster documentation (covering specifications for clusters, commands, and attributes). query (str, req); top_k (int, req).
schedule_workflow Schedule a sequential workflow of steps at a virtual absolute time. The scheduled time must be in the future relative to the current time. start_time (str, req; “YYYY-$M$-$D$ $H$:$M$:SS”); steps (list, req; e.g., {"tool":…, "args":…}).
get_current_time Get current virtual time as human-friendly string “YYYY-$M$-$D$ $H$:$M$:SS”. (none)
get_workflow_list Get list of workflows with optional filtering. (none)
A.3 List of Matter Clusters
Table 4: Implemented Matter clusters.
Cluster Attributes Commands
Basic Information VendorName, VendorID, ProductName, ProductID None
Descriptor DeviceTypeList, ServerList, ClientList, PartsList, TagList None
OnOff GlobalSceneControl, OnTime, OffWaitTime, StartUpOnOff Off, On, Toggle
Level Control CurrentLevel, RemainingTime, MinLevel, MaxLevel, CurrentFrequency, MinFrequency, MaxFrequency, OnOffTransitionTime, OnLevel, OnTransitionTime, OffTransitionTime, DefaultMoveRate, Options, StartUpCurrentLevel MoveToLevel, Move, Step, Stop, MoveToClosestFrequency
Fan Control FanMode, FanModeSequence, PercentSetting, PercentCurrent Step
MediaPlayback CurrentState Play, Pause, Stop, StartOver, Previous, Next, Rewind, FastForward
Channel ChannelList, Lineup, CurrentChannel ChangeChannel, ChangeChannelByNumber, SkipChannel
KeypadInput SupportedKeys SendKey
Identify IdentifyTime, IdentifyType Identify, TriggerEffect
Operational State PhaseList, Current Phase, CountdownTime, Operational State List, Operational State, Operational Error Pause, Resume, Stop, Start, OperationalCommandResponse
Power Source ClusterRevision, FeatureMap, Status, Order, Description, EndpointList, WiredAssessedInputVoltage, BatVoltage, BatPercentRemaining, BatChargeState, ActiveBatFaults None
Power Topology ClusterRevision, FeatureMap, AvailableEndpoints, ActiveEndpoints None
Electrical Power Measurement PowerMode, NumberOfMeasurementTypes, Accuracy, ReactiveCurrent, ApparentCurrent, ReactivePower, ApparentPower, RMSVoltage, RMSCurrent, RMSPower, Frequency, PowerFactor StartMeasurement, StopMeasurement, ResetMeasurement, GetMeasurementSnapshot
Electrical Energy Measurement Accuracy, CumulativeEnergyImported, CumulativeEnergyExported, PeriodicEnergyImported, PeriodicEnergyExported, CumulativeEnergyReset StartEnergyMeasurement, StopEnergyMeasurement, ResetCumulativeEnergy, GetEnergySnapshot
Device Energy Management ESAType, ESACanGenerate, ESAState, AbsMinPower, AbsMaxPower, PowerAdjustmentCapability, Forecast, OptOutState None
Dishwasher Mode SupportedModes, CurrentMode ChangeToMode, GetSupportedModes
Dishwasher Alarm Mask, Latch, State, Supported Reset, ModifyEnabledAlarms, GetAlarmState, GetActiveAlarms
Refrigerator And Temperature Controlled Cabinet Mode SupportedModes, CurrentMode ChangeToMode
RVC Clean Mode SupportedModes, CurrentMode ChangeToMode
RVC Operational State PhaseList, CurrentPhase, CountdownTime, Operational StateList, Operational State, OperationalError Pause, Resume, GoHome
RVC Run Mode SupportedModes, CurrentMode Start, Stop, Map, StopMap
Temperature Control TemperatureSetpoint, MinTemperature, MaxTemperature, Step, SelectedTemperatureLevel, SupportedTemperatureLevels SetTemperature
Temperature Measurement MeasuredValue, MinMeasuredValue, MaxMeasuredValue None
Thermostat LocalTemperature, OccupiedCoolingSetpoint, OccupiedHeatingSetpoint, ControlSequenceOfOperation, SystemMode SetpointRaiseLower
WindowCovering Type, ConfigStatus, OperationalStatus, EndProductType, Mode, SafetyStatus, CurrentPositionLiftPercent100ths, TargetPositionLiftPercent100ths, NumberOfActuationsLift, etc. UpOrOpen, DownOrClose, StopMotion, GoToLiftPercentage
Laundry Dryer Controls SupportedDrynessLevels, SelectedDrynessLevel None
Laundry Dryer Mode SupportedModes, CurrentMode ChangeToMode
Laundry Washer Controls SpinSpeeds, SpinSpeedCurrent, NumberOfRinses, SupportedRinses None
Laundry Washer Mode SupportedModes ChangeToMode
Relative Humidity Measurement MeasuredValue, MinMeasuredValue, MaxMeasuredValue, Tolerance None
A.4 List of Device Types
Table 5: List of implemented device types and their corresponding clusters.
Device type Clusters
Air Conditioner Basic Information, Fan Control, OnOff, Thermostat
Air Purifier Basic Information, Descriptor, Fan Control, Identify, OnOff
Dehumidifier Basic Information, Fan Control, OnOff, Relative Humidity Measurement
Dimmable Light Basic Information, Level Control, OnOff
Dishwasher Basic Information, OnOff, Operational State
Electrical Sensor Basic Information, Electrical Energy Measurement, Electrical Power Measurement, Power Topology
Fan Basic Information, Fan Control, OnOff
Freezer Basic Information, Descriptor, Refrigerator And Temperature Controlled Cabinet Mode, Temperature Control, Temperature Measurement
Heat Pump Basic Information, Descriptor, Device Energy Management, Electrical Energy Measurement, Electrical Power Measurement, Power Source, Power Topology, Thermostat
Humidifier Basic Information, Fan Control, OnOff, Relative Humidity Measurement
Laundry Dryer Basic Information, Laundry Dryer Controls, Laundry Dryer Mode, OnOff, Operational State
Laundry Washer Basic Information, Laundry Washer Controls, LaundryWasherMode, OnOff, Operational State, Temperature Control
On Off Light Basic Information, OnOff
Refrigerator Basic Information, Descriptor, Refrigerator And Temperature Controlled Cabinet Mode, Temperature Control, Temperature Measurement
RVC Basic Information, RVCCleanMode, RVCOperational State, RVCRunMode
TV Basic Information, Channel, KeypadInput, Level Control, MediaPlayback, OnOff
Window Covering Controller Basic Information, Window Covering
A.5 Goal Examples
Table 6: Example goals for each query types
Query Type Query Required Actions Goal
QT1 Feasible How bright is the utility room lighting right now? I am sorting some boxes and wondering if there is enough light. Also how is the living room humidity doing? I am thinking about the plants there and want to know if they are comfortable. get_room_states(utility_room) get_room_states(living_room) The utility room illuminance is 1000 lux. The living room humidity is 50%.
QT1 Infeasible I am about to shower and wondering what fan modes are available for fan 1 in the bathroom? get_room_devices(bathroom) Bathroom fan 1 not found; mode unavailable.
QT2 Feasible Ugh the kitchen feels really dry my hands are tight I left the bread rising there so I am already a bit worried about it. The living room feels dusty my eyes are itching and my throat is a little raw like there is grit in the air. get_room_devices(kitchen) get_room_devices(living_room) Increase kitchen humidity; decrease living room PM10.
QT2 Infeasible Ugh the office is so chilly, my hands go numb just thinking about working there later get_room_devices(office) Office heat pump 1 is missing; cannot increase temperature.
QT3 Feasible Set a softer light in the living room for evening reading, turn the living room dimmer light 1 on and set it to level 50. Cool the study a bit for working comfort, turn the study room AC 1 on, switch it to cooling mode and set the fan to 50 percent. get_room_devices(living_room) get_room_devices(study_room) Living room dimmable light 1 on at level 50; study room air conditioner 1 on, cooling mode, fan 50%.
QT3 Infeasible It’s a bit stuffy this morning, please turn on the bedroom air purifier 1 and set the fan to 80 percent. get_room_devices(bedroom) Not feasible: bedroom air purifier 1 is missing; cannot set fan to 80%.
QT4-1 While I am out here sorting laundry and trying to clear damp air, get the bathroom comfortable so it feels fresh by the time I walk over. Power on fan 1 in the bathroom 9 minutes from now at 30 percent, and bump it up to 40 percent 7 minutes after the prior action. Power on dimmer light 1 in the bathroom 28 minutes from now at level 10, and raise it to level 40 17 minutes after the prior action. get_room_devices(bathroom) At 9 minutes: bathroom fan 1 on, 30%. At 16 minutes: bathroom fan 1 on, 40%. At 28 minutes: bathroom dimmable light 1 on, level 10. At 45 minutes: bathroom dimmable light 1 on, level 40.
QT4-1 Temporal-Conflict Can you from the kitchen schedule dimmer light 1 in the living room to turn on and set to 80 percent in eight minutes from now, which will be 11:25 AM, I need it like that to warm up the room for guests and the start of the movie None At 8 minutes: living room dimmable light 1 on, level 80.
QT4-2 I am folding laundry and getting things ready. 20 minutes after the washer 1 in the utility room finishes, power on air purifier 1 in the living room and set the fan to 40 percent and switch heat pump 1 in the utility room to heating mode get_room_devices(living_room) get_room_devices(utility_room) At 79 minutes: living room air purifier 1 on, fan 40%; utility room heat pump 1 in heating mode.
QT4-2 Temporal-Conflict The wash leaves the utility room humid and cool so I want the air cleaned and the space warmed right after it settles. Exactly 20 minutes after washer 1 in the utility room finishes and at 12 36 PM, turn on air purifier 1 in the living room to a gentle fan speed and turn on heat pump 1 in the utility room for heating. None At 79 minutes: living room air purifier 1 on, fan 40%; utility room heat pump 1 in heating mode.
QT4-3 Waiting on the kitchen steam to clear so the laundry does not get musty. When dishwasher 1 in the kitchen finishes wait 11 minutes. Then start dryer 1 in the utility room. Set it to running and dryness level 1. get_room_devices(utility_room) At 99 minutes: utility room dryer 1 stopped. At 100 minutes: utility room dryer 1 running with dryness level 1.
QT4-3 Temporal-Conflict Start dryer 1 in the bathroom at twelve thirty six PM. Pause dryer 1 in the bathroom immediately when dryer 1 in the utility room finishes to avoid tripping the breaker and keep the laundry loads in order. None At 43 minutes: bathroom dryer 1 running with dryness level 1; at 44 minutes: bathroom dryer 1 paused.
A.6 LLM Judge Validation
To validate the LLM-based judging, we compared its assessments to human labels on a random subset of 70 episodes spanning all judge-scored tasks. Human annotators showed very high inter-rater reliability (Cohen’s $\kappa$ = 0.913). The LLM-Judge achieved substantial agreement with the consensus human labels (Cohen’s $\kappa$ = 0.826). These results support using the LLM-Judge as a reliable substitute for human evaluation in our benchmark.
After manually reviewing the 155 cases that the LLM-Judge evaluated as incorrect, we found that only 5 were misclassifications, underscoring the reliability of the evaluation. The detailed error distributions can be found in Table  A.11.
A.7 Experimental Setup
We evaluate 11 recent closed and open-source models: Llama-4-Scout (Meta AI, 2024), Llama-4-Maverick (Meta AI, 2024), Qwen3-32B (Yang et al., 2025), Qwen3-235B-A22B (Yang et al., 2025), Gemma-3-12B-it (Team et al., 2025), Gemma-3-27B-it (Team et al., 2025), Gemini-2.5-Flash (Comanici et al., 2025), Gemini-2.5-Flash-Lite (Comanici et al., 2025), GPT-4.1, GPT-4.1-mini (OpenAI, 2025a), and GPT-4.1-nano (OpenAI, 2025a). All models were accessed via the OpenRouter API (OpenRouter, 2025) to ensure standardized access and comparability.
A.8 ReAct Prompt
ReAct Prompt
A.9 LLM Judge Prompt
This section presents examples of the LLM Judge Prompts that were used to evaluate smart home LLM agents. Each query type (QT) has a dedicated evaluation prompt with specific criteria.
A.9.1 QT1 Feasible Judge Prompt
QT1 Feasible Judge Prompt (Normal)
A.9.2 QT1 Infeasible Judge Prompt
QT1 Infeasible Judge Prompt
A.9.3 QT2 Infeasible Judge Prompt
QT2 Infeasible Judge Prompt
A.9.4 QT2 Infeasible-Nonexistence Judge Prompt
QT2 Infeasible Judge Prompt
A.9.5 QT3 Infeasible Judge Prompt
QT3 Infeasible Judge Prompt
A.9.6 QT4-1 Judge Prompt
QT4-1 Judge Prompt
A.9.7 QT4-2 Judge Prompt
QT4-2 Judge Prompt
A.9.8 QT4-3 Judge Prompt
QT4-3 Judge Prompt
A.10 Error Taxonomy Details
Table 7: Error Types in Feasible Episodes
Error Type Definition Example
Environment Perception Errors (EP) Failure to correctly perceive or retrieve a value of environmental variables. Querying wrong sensor, misidentifying device state, guessing instead of perceiving.
Intent Inference Errors (II) Misinterpreting user’s underlying goal. Not executing actual commands even when a user’s intention is clear.
Device Control Errors (DC) Executing the wrong device, wrong command, or missing control steps. setting wrong channel, adjusting fan speed without turning it on first.
Action Planning Errors (AP) Incorrect or incomplete construction of the control workflow. Breaking logical dependencies, only executing part of a multi-goal query without consideration.
Temporal Reasoning Errors (TR) Miscalculating relative/absolute times or sequence alignment. Scheduling “in 10 minutes” at wrong time, miscomputing dishwasher completion.
Table 8: Error Types in Infeasible Episodes
Error Type Definition Example
Contradiction Mishandling Errors (CM) The agent detects a contradiction but does not follow the proper instruction-following rule. e.g., instead of informing the user that a requested action is impossible, it arbitrarily manipulates other devices or ignores the instruction.
Contradiction Blindness Errors (CB) The agent completely fails to recognize a contradiction and executes the request as if it were valid. e.g., dimming an on/off light, scheduling conflicting temporal actions without noticing inconsistency.
LLM-Judge Errors (LJ) Errors caused not by the agent but by the evaluation system misclassifying or overlooking behavior. e.g., penalizing an informative refusal as a failure, or wrongly accepting hallucinated control as valid.
A.11 Error Type Distributions
Error Type QT2 QT3 QT4-1 QT4-2 QT4-3
Environment Perception (EP) 3 0 4 1 0
Intent Inference (II) 3 1 0 4 5
Device Control (DC) 20 7 13 13 8
Action Planning (AP) 2 0 6 3 7
Temporal Reasoning (TR) 0 0 2 6 13
Total 28 8 25 27 33
Table 9: Error type distribution of GPT-4.1 in feasible episodes.
Error Types QT1 QT2 QT3 QT4-1 QT4-2 QT4-3
Contradiction Mishandling (CM) 8 24 6 5 6 1
Contradiction Blindness (CB) 0 5 0 40 25 30
LLM-Judge (LJ) 1 0 0 0 1 2
Total 9 29 6 45 32 33
Table 10: Error type distribution of GPT-4.1 in infeasible episodes.
A.12 Equations of Aggregators
$S_{r,t+1}=S_{r,t}+\sum_{d\in D_{S,r}}\Delta S_{d,r}(t),$ (1)
where $D_{S,r}$ denotes the set of devices in room $r$ that are defined to affect state $S$, and $\Delta S_{d,r}(t)$ represents the contribution of device $d$ at tick $t$ to $S$ in room $r$.
A.13 QT3 Error Type Distribution
Figure 6: Distribution of tool execution error types in successful QT3 Feasible cases across four models.
A.14 LLM Usage
We disclose that LLMs (OpenAI GPT-5) were used as a general-purpose assistant for language editing, including polishing wording, improving clarity, and maintaining consistency. LLMs were not involved in research design, implementation, experiments, or analysis. All scientific contributions and claims in this paper are the sole responsibility of the authors.

### v2

@@
 To address this, we introduce SimuHome, a time-accelerated home environment that simulates smart devices, supports API calls, and reflects changes in environmental variables.
-By building the simulator on the Matter protocol [^1] https://csa-iot.org/all-solutions/matter/—the global industry standard for smart home communication—SimuHome provides a high-fidelity environment, and agents validated in SimuHome can be deployed on real Matter-compliant devices with minimal adaptation.
+By building the simulator on the Matter protocol [^1] https://csa-iot.org/all-solutions/matter/, the global industry standard for smart home communication, SimuHome provides a high-fidelity environment, and agents validated in SimuHome can be deployed on real Matter-compliant devices with minimal adaptation.
 We provide a challenging benchmark of 600 episodes across twelve user query types that require the aforementioned capabilities.
-Our evaluation of 11 agents under a unified ReAct framework reveals that while models perform well on simple tasks, they struggle with latent intent inference, state verification, and especially temporal scheduling.
-Even the top-performing model, GPT-4.1, reaches only 54% success rate.
-These findings highlight a critical need for methods that can reliably verify the current state via tools before acting and coordinate time-dependent actions.
+Our evaluation of 16 agents under a unified ReAct framework reveals distinct capabilities and limitations across models.
+Models under 7B parameters exhibited negligible performance across all query types.
+Even GPT-4.1, the best-performing standard model, struggled with implicit intent inference, state verification, and particularly temporal scheduling.
+While reasoning models such as GPT-5.1 consistently outperformed standard models on every query type, they required over three times the average inference time, which can be prohibitive for real-time smart home applications.
+This highlights a critical trade-off between task performance and real-world practicality.
 [^2] We will release our code and dataset upon publication of the paper.
@@
 Infeasible cases, which embed false premises, physical limits, or temporal conflicts, are assessed by LLM judges.
-We evaluate 11 LLM agents under a unified ReAct (Yao et al., 2023) setup across 600 episodes with feasible and infeasible variants, scoring feasible tasks by simulator state comparisons and assessing infeasible logic checks with validated LLM judges.
-Models handle simple retrieval and explicit device control comparatively well.
-They often fail to infer latent intent or to verify current device and environment states before acting.
-Temporal scheduling is the most challenging area: contradiction blindness and mishandling are common, and even GPT-4.1 reaches only 54% success rate.
-These results motivate methods that verify the current state via tools before acting and that plan and coordinate time-dependent actions reliably in dynamic home environments.
+We evaluate 16 LLM agents under a unified ReAct (Yao et al., 2023) setup across 600 episodes with feasible and infeasible variants, scoring feasible tasks by simulator state comparisons and assessing infeasible tasks with validated LLM judges.
+Standard models handle simple retrieval and explicit device control well, but struggle to infer latent intent and verify current states before acting.
+Among all query types, temporal scheduling proves most challenging.
+Even GPT-4.1, the best-performing standard model, achieves only 12–50% success.
+Reasoning models such as GPT-5.1 improve substantially (44–100%), but their threefold inference overhead limits their use in real-time applications.
+This motivates developing efficient methods that verify system state via tools and reliably coordinate time-dependent actions.
 2 Related Work
@@
 In total, we model 17 distinct device types.
-A full list of these devices can be found in Appendix A.4.
+A full list of these devices can be found in Appendix LABEL:app:device_list.
 Real-Time State Update Mechanism.
@@
 This mechanism ensures that the environment responds realistically to agent actions.
-The detailed update equations for the Aggregator are provided in Appendix A.12.
+The detailed update equations for the Aggregator are provided in Appendix LABEL:app:aggregator_equation.
 Agent-Simulator Interface.
-The agent interacts with the simulator by invoking a set of 13 tools.
+The agent interacts with the simulator by invoking a set of 17 tools.
 The structure of these tools mirrors Matter’s modular approach to defining device capabilities.
-Detailed tool specifications are provided in Appendix A.2.
+Detailed tool specifications are provided in Appendix B.
 3.3 Task Definition
@@
 Each type is paired with an infeasible scenario to test the agent’s capacity for logical consistency and constraint handling, yielding a total of 12 categories.
-See Appendix A.1 for examples of infeasible scenarios corresponding to each query type.
+See Appendix A for examples of infeasible scenarios corresponding to each query type.
 QT1 (Environment Perception).
@@
 Figure 2: Episode Generation Pipeline
-Figure 3: Episode Evaluation Pipeline
 Definition and Components of Episode.
@@
 A goal defines the desired final state of specific devices or environmental variables that the agent must achieve.
-The generation process, which varies by query type (see Appendix A.5), is designed to ensure all goals are logically consistent.
+The generation process, which varies by query type (see Appendix LABEL:app:goal_examples), is designed to ensure all goals are logically consistent.
 For instance, as illustrated for QT3 in Figure 2 (Step 3), a device goal is created by sampling from a pre-defined set of valid states (e.g., onoff: on, fanspeed: 40%).
@@
 Therefore, we first defined a verifiable goal for the agent to accomplish and subsequently synthesized a natural language query based on it.
-We then used GPT-5-mini (OpenAI, 2025b) to synthesize the natural language queries from these predefined goals.
+We then used GPT-5-mini (OpenAI, 2025c) to synthesize the natural language queries from these predefined goals.
 To ensure each query accurately reflected its predefined goal, two graduate students researching tool agents independently reviewed the entire dataset.
@@
 We generated 50 distinct episodes for each of the 12 query types, resulting in a high-quality dataset of 600 episodes designed for evaluating smart home agents.
-5 Evaluation
+Figure 3: Episode Evaluation Pipeline
 4.3 Evaluation Methods
@@
 In the QT4-Feasible episode shown in Figure 3, after the agent completes all its actions, the simulator accelerates time to 19:00, when the laundry cycle finishes.
-Subsequently, the goal state—the washer being off and the kitchen light being on—is compared with the simulator’s final state to determine success.
-This direct state comparison yields a clear and fully automated success criterion and enables fair model-to-model comparisons under the same conditions.
+The goal state (washer off, kitchen light on) is then compared with the simulator’s final state to determine success.
+This direct state comparison enables fully automated and fair model-to-model comparisons.
+We apply simulator-based evaluation to all feasible episodes in QT2, QT3, and QT4, which involve physical state changes in the home environment.
 LLM-judge-based Evaluation.
-We employ an LLM-based judge for episodes where success depends not on physical state changes, but on the agent’s final natural-language response.
-This provides an objective and effective complement to simulator-based evaluation.
-In our procedure, the judge is provided with the episode goal and the user query, along with a concise description of any infeasible conditions that must be verified.
-We also provide the agent’s full ReAct trajectory so that the judge can verify whether the final answer is supported by a coherent and sufficient chain of reasoning.
-For example, in the QT4-Infeasible episode shown in Figure 3, the LLM-judge assesses the agent’s explanation of a scheduling conflict.
-The judge scores only the final user-facing answer while using the trajectory to confirm consistency.
-For each case, we asked the judge three times and adopted the consistent outcome to ensure reliability (Taubenfeld et al., 2025).
-Detailed prompt templates and rubrics are available in Appendix A.9.
-To validate our LLM-judges, we confirmed that their assessments achieved substantial agreement (Cohen’s $\kappa$ = 0.826) with human evaluations, which themselves showed very high inter-rater reliability (see Appendix A.6 for the full analysis).
-5.2 Experiments
-Experimental Setup.
-We quantitatively compare models’ inference, planning, and tool-calling performance across the 12 query types defined in §4.1.
-The evaluation covers 11 models spanning open-source (Llama, Qwen, Gemma) and closed-source (Gemini, GPT) families, with reproducibility details provided in Appendix A.7.
-We employ the ReAct framework (Yao et al., 2023) for all experiments, enabling the models to generate both reasoning traces and actions in a step-by-step manner within our smart home simulation environment.
-The specific agent prompt used for the ReAct framework is provided in Appendix A.8.
-5.3 Main Results
+We employ an LLM-based judge for episodes where success depends on the agent’s final natural-language response rather than physical state changes.
+The judge receives the episode goal, user query, and the agent’s full reasoning trajectory, along with a description of any infeasible conditions that must be verified.
+This allows the judge to assess whether the final answer is supported by coherent reasoning.
+For example, in the QT4-Infeasible episode shown in Figure 3, the LLM-judge evaluates the agent’s explanation of a scheduling conflict.
+We apply LLM-judge-based evaluation to all infeasible episodes (QT1-IF through QT4-IF), which require assessing whether the agent correctly identifies and explains constraint violations.
+Additionally, QT1-Feasible also uses LLM-judge evaluation because success depends on providing accurate information in natural language rather than changing device states.
+For reliability, we query the judge three times per case and adopt the consistent outcome (Taubenfeld et al., 2025).
+Our LLM-judges achieved substantial agreement (Cohen’s $\kappa$ = 0.826) with human evaluations (see Appendix LABEL:app:llm_judge_validation).
+Detailed prompt templates are in Appendix LABEL:app:llm_judge_prompt.
+5 Experiments
 Table 1: Evaluation results show success rates (in %) across query types (QTs).
@@
 QT1 QT2 QT3 QT4-1 QT4-2 QT4-3
-Model FJ IFJ FS IFJ FS IFJ FS IFJ FS IFJ FS IFJ
-Llama-4-Scout $58$ $42$ 22 $22$ $24$ $34$ 44 44 22 22 22 0
-Llama-4-Maverick 96 $78$ $52$ $36$ $88$ $74$ $22$ $14$ $18$ $10$ 32 88
-Qwen3-32B $82$ $66$ 62 $30$ $52$ $68$ $18$ $14$ $14$ 88 $16$ 66
+Models FJ IFJ FS IFJ FS IFJ FS IFJ FS IFJ FS IFJ
+Open Source Large Language Models (<7B)
+Llama3.2-1B-it 0 0 0 0 0 0 0 0 0 0 0 0
+Llama3.2-3B-it 10 12 0 2 4 0 2 0 2 0 0 0
+Gemma3-4B-it 44 32 12 10 28 8 0 0 2 0 0 4
+Open Source Large Language Models (Standard)
+Llama4-Scout 58 42 2 22 24 34 4 4 2 2 2 0
+Llama4-Maverick 96 78 52 36 88 74 22 14 18 10 32 8
+Qwen3-32B 82 66 62 30 52 68 18 14 14 8 16 6
 Qwen3-235B-A22B 86 74 32 36 84 70 26 18 38 34 28 48
-Gemma-3-12B-it $78$ $38$ $14$ $32$ $32$ $24$ 22 0 0 0 0 0
-Gemma-3-27B-it $80$ $48$ $54$ $24$ $48$ $44$ 44 22 $10$ 88 0 66
-Gemini-2.5-Flash-Lite $78$ $60$ $44$ 50 $50$ $50$ 88 $34$ $10$ $16$ $16$ 2020
-Gemini-2.5-Flash $92$ $86$ $66$ $54$ $82$ $74$ $22$ $44$ 40 32 $12$ 32
-GPT-4.1-nano $58$ $42$ 66 $12$ $30$ $16$ 22 66 66 0 0 0
-GPT-4.1-mini 96 $76$ 62 $28$ $64$ 76 26 40 40 2020 $10$ $28$
+Gemma3-12B-it 78 38 14 32 32 24 2 0 0 0 0 0
+Gemma3-27B-it 80 48 54 24 48 44 4 2 10 8 0 6
+Closed Source Large Language Models (Standard)
+Gemini2.5-Flash-Lite 78 60 44 50 50 50 8 34 10 16 16 20
+Gemini2.5-Flash 92 86 66 54 82 74 22 44 40 32 12 32
+GPT-4.1-nano 58 42 6 12 30 16 2 6 6 0 0 0
+GPT-4.1-mini 96 76 62 28 64 76 26 40 40 20 10 28
 GPT-4.1 98 82 44 44 84 88 50 12 46 34 34 32
-Table 1 reports performance across all query types (QT1-QT4) under both feasible (F) and infeasible (IF) conditions. Several consistent patterns emerge.
-Environment perception (QT1).
-QT1-F episodes are largely solved: Most frontier models exceeded 85%.
-In QT1-IF, we observe a success rate degradation across models, reflecting a consistent challenge in detecting requests based on false premises, such as references to non-existent devices.
-Explicit vs.
-implicit device control (QT2 vs.
-QT3).
-Across both feasible and infeasible settings, most models handle explicit commands (QT3) substantially better than implicit requests (QT2).
-In feasible episodes, leading models exceed 84% on QT3, whereas QT2 plateaus at 62-66%.
-In feasible settings, models consistently perform better when users issue explicit commands (QT3) compared to implicit requests (QT2).
-In infeasible settings, however, the contrast becomes more pronounced: models can often detect nonexistent devices in QT3-IF with relatively high success rate, but they struggle far more when the device exists yet the goal is unattainable due to operational constraints.
-For example, an air conditioner already operating at maximum fan speed cannot lower the temperature further.
-This indicates that reasoning over such constraint-driven infeasibility poses a greater challenge for current models.
-Temporal reasoning (QT4).
-QT4 queries, which require agents to handle scheduling, coordination, and temporal consistency, were the most challenging across both feasible and infeasible settings.
-GPT-4.1 achieved 50% on QT4-1, 46% on QT4-2, and 34% on QT4-3, outperforming peers but still far from robust.
-In infeasible cases with contradictory or impossible timing, performance degraded sharply: even GPT-4.1 and Gemini-2.5-Flash peaked at 44% (QT4-1-IF) and 34% (QT4-2-IF), with Qwen3-235B-A22B delivering notable performance at 48% on QT4-3-IF.
-These results highlight that contradiction detection and temporal validation remain open challenges.
-Model-level comparison.
-Closed-source models (GPT-4.1, Gemini-2.5-Flash) consistently led overall, but none achieved stable performance across both feasible and infeasible tasks.
-Mid-sized open-source models such as Qwen3-235B-A22B and Gemini-2.5-Flash-Lite produced notable results on certain query types—for instance, Qwen3-235B-A22B reached 48% on QT4-3-IF—but their performance was uneven across query types.
-Smaller models (Gemma-3-12B-it, GPT-4.1-nano) remained limited overall, particularly on temporally demanding tasks.
-Table 2: Error taxonomy. Detailed descriptions and examples are provided in Appendix A.10.
+Closed Source Large Language Models (with Reasoning)
+Gemini2.5-Pro 96 78 60 56 76 72 44 94 60 76 46 50
+GPT-5.1 100 94 80 50 86 92 60 100 72 92 56 44
+Experimental Setup.
+We evaluate 16 models across the 12 query types defined in §4.1, spanning four categories: open-source models under 7B parameters, open-source standard models, closed-source standard models, and closed-source reasoning models.
+All experiments use the ReAct framework (Yao et al., 2023), enabling step-by-step reasoning and action generation.
+Reproducibility details and agent prompts are in Appendix LABEL:app:experimental_setup and LABEL:app:react_prompt.
+5.1 Main Results
+Table 1 presents the performance across all query types (QT1-QT4). We analyze the results by categorizing models into three groups based on scale and reasoning capability.
+Lightweight Models.
+Models under 7B exhibit severe limitations across all query types.
+For instance, Llama3.2-1B-it achieves 0% across all tasks, while Gemma3-4B-it shows marginal success on basic tasks such as information retrieval (QT1-F: 44%) and explicit commands (QT3-F: 28%).
+However, these models completely fail on complex reasoning tasks, with near-zero success rates on implicit intent inference (QT2) and temporal reasoning (QT4), indicating insufficient reasoning capabilities for our benchmark.
+Standard Models.
+This group comprises closed-source models (e.g., GPT-4.1, Gemini2.5-Flash) and large open-source models (e.g., Llama4-Maverick).
+Even GPT-4.1, the best-performing standard model, struggles with implicit intent inference (QT2) and temporal reasoning (QT4).
+GPT-4.1 achieves only 44% on QT2-F, compared to 84% on explicit device control (QT3-F).
+This gap highlights the difficulty of interpreting ambiguous user queries.
+Performance on temporal reasoning (QT4) plateaus at approximately 30–50% for feasible episodes and degrades further in infeasible scenarios (12–34%), revealing limited ability to detect temporal contradictions.
+Reasoning Models.
+Advanced reasoning models (Gemini2.5-Pro, GPT-5.1) demonstrate substantial improvements over standard models, particularly in temporal reasoning and implicit intent inference.
+On temporal reasoning tasks (QT4-F), GPT-5.1 achieves 56–72% compared to GPT-4.1’s 34–50%.
+The improvement is especially pronounced in infeasible episodes, where GPT-5.1 reaches 100% on QT4-1-IF and 92% on QT4-2-IF, compared to GPT-4.1’s 12% and 34%.
+This demonstrates stronger capability in detecting temporal contradictions.
+Beyond temporal reasoning, GPT-5.1 also achieves 80% on QT2-F compared to GPT-4.1’s 44%, indicating improved interpretation of latent user intents.
+5.2 Analysis
+Table 2: Error taxonomy. Detailed descriptions and examples are provided in Appendix LABEL:app:error_taxonomy_details.
 Category Error Type Definition
@@
 LLM-Judge (LJ) Misclassification by LLM-Judge.
-Figure 4: Error type distributions of GPT-4.1 on QT2 and QT4. (a) and (b) show the distributions for feasible episodes, while (c) and (d) present the distributions for infeasible episodes.
-6 Analysis
-6.1 Error Analysis
+Figure 4: Error type distributions of GPT-4.1 on QT2 and QT4.
+5.2.1 Error Analysis.
 We define eight error types to analyze agent failures: five for feasible episodes (EP, II, DC, AP, TR) and three for infeasible episodes (CM, CB, LJ).
@@
 Even when contradictions were recognized, responses were frequently mishandled (CM).
-6.2 Role of Tool Feedback
+Table 3: Average episode completion time (seconds) across query types. While reasoning models achieve high accuracy, they incur significant latency costs compared to the standard model (GPT-4.1).
+QT1 QT2 QT3 QT4-1 QT4-2 QT4-3
+Model F IF F IF F IF F IF F IF F IF
+Gemini-2.5-Pro 24.1 22.4 57.5 48.8 66.1 27.8 74.0 12.5 57.7 37.0 53.7 53.1
+GPT-5.1 35.7 38.4 109.4 99.6 78.6 54.3 121.1 13.5 135.1 76.0 112.7 111.0
+GPT-4.1 8.3 7.8 23.6 20.2 22.9 9.4 26.6 12.3 28.7 23.7 29.7 25.9
 Figure 5: Tool-call error patterns of four models on QT3-F.
@@
 The right chart shows the proportion of tasks achieved through first-try success versus those requiring error recovery.
+5.2.2 Role of Tool Feedback
 To better understand agent dynamics, we examined QT3, where most models were relatively strong.
@@
 Consequently, the simulator provides little corrective signal, leaving the agent unable to revise its plan.
+5.2.3 Performance-Latency Trade-off
+As discussed in §5.1, advanced reasoning models demonstrate significant performance gains.
+However, these gains come with a critical trade-off in latency.
+Table 3 shows the average episode completion time for high-performing models.
+Reasoning models have unacceptably large latency for practical deployment.
+For instance, GPT-5.1 takes up to 135.1 seconds and Gemini-2.5-Pro takes up to 74.0 seconds on average.
+In a smart home context, users typically expect immediate responses in less than a second.
+Therefore, even GPT-4.1 (best-performing standard model) cannot meet real-time requirements, as it takes up to 29.7 seconds.
+The extended latency of reasoning models further limits practical utility despite their accuracy improvements.
+At the other end of the spectrum, models under 7B parameters are more practical choices for deployment but achieved minimal performance as shown in §5.1.
+This suggests that the complex reasoning and tool-use coordination required by SimuHome present significant challenges for real-world smart home environments.
+5.2.4 Disentangling Framework Limitations from Model Capabilities
+Figure 6: Performance comparison between ReAct and HiAgent on QT4 tasks.
+Table 4: Recovery rates from QT4 failures using self-review.
+Type Failures Recoveries Rate Steps
+QT4-1 25 2 8.0% 64.4
+QT4-2 27 5 18.5% 26.8
+QT4-3 33 0 0.0% 29.7
+To investigate whether the failures in QT4 stem from the ReAct framework itself or from the models’ core reasoning capabilities, we conducted two ablation studies.
+Adopting a Complex Planning Algorithm.
+We adopted HiAgent (Hu et al., 2025), a framework that incorporates hierarchical working memory and advanced planning algorithms.
+We compared its performance against ReAct on QT4 tasks using GPT-4.1.
+Figure 6 shows that HiAgent outperformed ReAct on QT4-2 and QT4-3, but underperformed on QT4-1, achieving only 40% compared to ReAct’s 50%.
+These results indicate that while ReAct’s framework has limited support for complex temporal coordination, not all temporal reasoning failures stem from the framework.
+Enabling Runtime Periodic Self-Review.
+We tested a realistic setting where agents can adjust plans through periodic self-review.
+After scheduling workflows, the agent received callback triggers to review and correct its plans before and immediately after execution.
+The agent was required to independently evaluate the outcome to determine the appropriate subsequent actions.
+The results in Table 4 reveal that self-review achieved only 0% to 18.5% recovery rates across QT4 tasks.
+Crucially, even when the agent was explicitly prompted to inspect the home state immediately after execution, it failed to recognize the task failure in most cases.
+The continued failure on QT4 tasks, even after adopting advanced planning techniques and enabling runtime periodic self-review, demonstrates that the primary bottleneck is not the ReAct framework itself.
+Instead, the failures are likely attributed to the inherently limited capabilities of current models to ground complex home states and perform precise temporal reasoning.
 6 Conclusion
@@
 We score feasible tasks by final state-to-goal comparison in the simulator and assess infeasible logic checks with LLM judge that shows high agreement with human evaluation.
-We evaluate 11 agents under the ReAct setup.
-Current open- and closed-source LLMs handle explicit control and simple retrieval well but often fail at latent-intent inference, live-state verification, and temporal scheduling; GPT-4.1 achieves an overall success rate of 54% across all query types.
-Appendix A Appendix
-A.1 Infeasible Query Types
+We evaluate 16 models under the ReAct setup.
+Even high-performing standard models struggle with implicit intent inference, state verification, and temporal scheduling.
+While reasoning models consistently outperform standard models across all query types, they incur prohibitive latency costs.
+These results highlight a critical performance-practicality trade-off, positioning SimuHome as a vital testbed to address these challenges for real-world deployment.
+Appendix A Infeasible Query Types
 QT1 Infeasible.
@@
 completion is not feasible and suggest the earliest possible completion time or an alternative sequential plan.
-A.2 List of Tools
-Table 3: Tool List for Agent
+Appendix B List of Tools
+Tool List for Agent
 Name Description Args
-finish Complete the task and return the final natural-language answer. answer (str, required): Final response text.
-execute_command Execute a command on a device (e.g., turn on light, set level, set setpoint).
-device_id (str, req); endpoint_id (int, req); cluster_id (str, req); command_id (str, req); args (dict, req).
-write_attribute Directly set a device attribute value. device_id (str, req); endpoint_id (int, req); cluster_id (str, req); attribute_id (str, req); value (any, req).
-get_all_attributes Get all attributes of a device. device_id (str, req).
-get_attribute Get a specific attribute of a device. device_id (str, req); endpoint_id (int, req); cluster_id (str, req); attribute_id (str, req).
-get_device_structure Get device structure (endpoints, clusters, attributes, and commands). device_id (str, req).
-get_rooms Get all rooms in the home along with their display names. (none)
-get_room_devices Get all devices in a room. room_id (str, req).
-get_room_states Get environmental states of a room (temperature, humidity, illuminance, PM10). room_id (str, req).
-get_cluster_doc Perform semantic search across Matter cluster documentation (covering specifications for clusters, commands, and attributes). query (str, req); top_k (int, req).
-schedule_workflow Schedule a sequential workflow of steps at a virtual absolute time.
-The scheduled time must be in the future relative to the current time.
-start_time (str, req; “YYYY-$M$-$D$ $H$:$M$:SS”); steps (list, req; e.g., {"tool":…, "args":…}).
-get_current_time Get current virtual time as human-friendly string “YYYY-$M$-$D$ $H$:$M$:SS”. (none)
-get_workflow_list Get list of workflows with optional filtering. (none)
-A.3 List of Matter Clusters
-Table 4: Implemented Matter clusters.
-Cluster Attributes Commands
-Basic Information VendorName, VendorID, ProductName, ProductID None
-Descriptor DeviceTypeList, ServerList, ClientList, PartsList, TagList None
-OnOff GlobalSceneControl, OnTime, OffWaitTime, StartUpOnOff Off, On, Toggle
-Level Control CurrentLevel, RemainingTime, MinLevel, MaxLevel, CurrentFrequency, MinFrequency, MaxFrequency, OnOffTransitionTime, OnLevel, OnTransitionTime, OffTransitionTime, DefaultMoveRate, Options, StartUpCurrentLevel MoveToLevel, Move, Step, Stop, MoveToClosestFrequency
-Fan Control FanMode, FanModeSequence, PercentSetting, PercentCurrent Step
-MediaPlayback CurrentState Play, Pause, Stop, StartOver, Previous, Next, Rewind, FastForward
-Channel ChannelList, Lineup, CurrentChannel ChangeChannel, ChangeChannelByNumber, SkipChannel
-KeypadInput SupportedKeys SendKey
-Identify IdentifyTime, IdentifyType Identify, TriggerEffect
-Operational State PhaseList, Current Phase, CountdownTime, Operational State List, Operational State, Operational Error Pause, Resume, Stop, Start, OperationalCommandResponse
-Power Source ClusterRevision, FeatureMap, Status, Order, Description, EndpointList, WiredAssessedInputVoltage, BatVoltage, BatPercentRemaining, BatChargeState, ActiveBatFaults None
-Power Topology ClusterRevision, FeatureMap, AvailableEndpoints, ActiveEndpoints None
-Electrical Power Measurement PowerMode, NumberOfMeasurementTypes, Accuracy, ReactiveCurrent, ApparentCurrent, ReactivePower, ApparentPower, RMSVoltage, RMSCurrent, RMSPower, Frequency, PowerFactor StartMeasurement, StopMeasurement, ResetMeasurement, GetMeasurementSnapshot
-Electrical Energy Measurement Accuracy, CumulativeEnergyImported, CumulativeEnergyExported, PeriodicEnergyImported, PeriodicEnergyExported, CumulativeEnergyReset StartEnergyMeasurement, StopEnergyMeasurement, ResetCumulativeEnergy, GetEnergySnapshot
-Device Energy Management ESAType, ESACanGenerate, ESAState, AbsMinPower, AbsMaxPower, PowerAdjustmentCapability, Forecast, OptOutState None
-Dishwasher Mode SupportedModes, CurrentMode ChangeToMode, GetSupportedModes
-Dishwasher Alarm Mask, Latch, State, Supported Reset, ModifyEnabledAlarms, GetAlarmState, GetActiveAlarms
-Refrigerator And Temperature Controlled Cabinet Mode SupportedModes, CurrentMode ChangeToMode
-RVC Clean Mode SupportedModes, CurrentMode ChangeToMode
-RVC Operational State PhaseList, CurrentPhase, CountdownTime, Operational StateList, Operational State, OperationalError Pause, Resume, GoHome
-RVC Run Mode SupportedModes, CurrentMode Start, Stop, Map, StopMap
-Temperature Control TemperatureSetpoint, MinTemperature, MaxTemperature, Step, SelectedTemperatureLevel, SupportedTemperatureLevels SetTemperature
-Temperature Measurement MeasuredValue, MinMeasuredValue, MaxMeasuredValue None
-Thermostat LocalTemperature, OccupiedCoolingSetpoint, OccupiedHeatingSetpoint, ControlSequenceOfOperation, SystemMode SetpointRaiseLower
-WindowCovering Type, ConfigStatus, OperationalStatus, EndProductType, Mode, SafetyStatus, CurrentPositionLiftPercent100ths, TargetPositionLiftPercent100ths, NumberOfActuationsLift, etc.
-UpOrOpen, DownOrClose, StopMotion, GoToLiftPercentage
-Laundry Dryer Controls SupportedDrynessLevels, SelectedDrynessLevel None
-Laundry Dryer Mode SupportedModes, CurrentMode ChangeToMode
-Laundry Washer Controls SpinSpeeds, SpinSpeedCurrent, NumberOfRinses, SupportedRinses None
-Laundry Washer Mode SupportedModes ChangeToMode
-Relative Humidity Measurement MeasuredValue, MinMeasuredValue, MaxMeasuredValue, Tolerance None
-A.4 List of Device Types
-Table 5: List of implemented device types and their corresponding clusters.
-Device type Clusters
-Air Conditioner Basic Information, Fan Control, OnOff, Thermostat
-Air Purifier Basic Information, Descriptor, Fan Control, Identify, OnOff
-Dehumidifier Basic Information, Fan Control, OnOff, Relative Humidity Measurement
-Dimmable Light Basic Information, Level Control, OnOff
-Dishwasher Basic Information, OnOff, Operational State
-Electrical Sensor Basic Information, Electrical Energy Measurement, Electrical Power Measurement, Power Topology
-Fan Basic Information, Fan Control, OnOff
-Freezer Basic Information, Descriptor, Refrigerator And Temperature Controlled Cabinet Mode, Temperature Control, Temperature Measurement
-Heat Pump Basic Information, Descriptor, Device Energy Management, Electrical Energy Measurement, Electrical Power Measurement, Power Source, Power Topology, Thermostat
-Humidifier Basic Information, Fan Control, OnOff, Relative Humidity Measurement
-Laundry Dryer Basic Information, Laundry Dryer Controls, Laundry Dryer Mode, OnOff, Operational State
-Laundry Washer Basic Information, Laundry Washer Controls, LaundryWasherMode, OnOff, Operational State, Temperature Control
-On Off Light Basic Information, OnOff
-Refrigerator Basic Information, Descriptor, Refrigerator And Temperature Controlled Cabinet Mode, Temperature Control, Temperature Measurement
-RVC Basic Information, RVCCleanMode, RVCOperational State, RVCRunMode
-TV Basic Information, Channel, KeypadInput, Level Control, MediaPlayback, OnOff
-Window Covering Controller Basic Information, Window Covering
-A.5 Goal Examples
-Table 6: Example goals for each query types
-Query Type Query Required Actions Goal
-QT1 Feasible How bright is the utility room lighting right now?
-I am sorting some boxes and wondering if there is enough light.
-Also how is the living room humidity doing?
-I am thinking about the plants there and want to know if they are comfortable.
-get_room_states(utility_room) get_room_states(living_room) The utility room illuminance is 1000 lux.
-The living room humidity is 50%.
-QT1 Infeasible I am about to shower and wondering what fan modes are available for fan 1 in the bathroom? get_room_devices(bathroom) Bathroom fan 1 not found; mode unavailable.
-QT2 Feasible Ugh the kitchen feels really dry my hands are tight I left the bread rising there so I am already a bit worried about it.
-The living room feels dusty my eyes are itching and my throat is a little raw like there is grit in the air.
-get_room_devices(kitchen) get_room_devices(living_room) Increase kitchen humidity; decrease living room PM10.
-QT2 Infeasible Ugh the office is so chilly, my hands go numb just thinking about working there later get_room_devices(office) Office heat pump 1 is missing; cannot increase temperature.
-QT3 Feasible Set a softer light in the living room for evening reading, turn the living room dimmer light 1 on and set it to level 50.
-Cool the study a bit for working comfort, turn the study room AC 1 on, switch it to cooling mode and set the fan to 50 percent.
-get_room_devices(living_room) get_room_devices(study_room) Living room dimmable light 1 on at level 50; study room air conditioner 1 on, cooling mode, fan 50%.
-QT3 Infeasible It’s a bit stuffy this morning, please turn on the bedroom air purifier 1 and set the fan to 80 percent.
-get_room_devices(bedroom) Not feasible: bedroom air purifier 1 is missing; cannot set fan to 80%.
-QT4-1 While I am out here sorting laundry and trying to clear damp air, get the bathroom comfortable so it feels fresh by the time I walk over.
-Power on fan 1 in the bathroom 9 minutes from now at 30 percent, and bump it up to 40 percent 7 minutes after the prior action.
-Power on dimmer light 1 in the bathroom 28 minutes from now at level 10, and raise it to level 40 17 minutes after the prior action.
-get_room_devices(bathroom) At 9 minutes: bathroom fan 1 on, 30%.
-At 16 minutes: bathroom fan 1 on, 40%.
-At 28 minutes: bathroom dimmable light 1 on, level 10.
-At 45 minutes: bathroom dimmable light 1 on, level 40.
-QT4-1 Temporal-Conflict Can you from the kitchen schedule dimmer light 1 in the living room to turn on and set to 80 percent in eight minutes from now, which will be 11:25 AM, I need it like that to warm up the room for guests and the start of the movie None At 8 minutes: living room dimmable light 1 on, level 80.
-QT4-2 I am folding laundry and getting things ready.
-20 minutes after the washer 1 in the utility room finishes, power on air purifier 1 in the living room and set the fan to 40 percent and switch heat pump 1 in the utility room to heating mode get_room_devices(living_room) get_room_devices(utility_room) At 79 minutes: living room air purifier 1 on, fan 40%; utility room heat pump 1 in heating mode.
-QT4-2 Temporal-Conflict The wash leaves the utility room humid and cool so I want the air cleaned and the space warmed right after it settles.
-Exactly 20 minutes after washer 1 in the utility room finishes and at 12 36 PM, turn on air purifier 1 in the living room to a gentle fan speed and turn on heat pump 1 in the utility room for heating.
-None At 79 minutes: living room air purifier 1 on, fan 40%; utility room heat pump 1 in heating mode.
-QT4-3 Waiting on the kitchen steam to clear so the laundry does not get musty.
-When dishwasher 1 in the kitchen finishes wait 11 minutes.
-Then start dryer 1 in the utility room.
-Set it to running and dryness level 1.
-get_room_devices(utility_room) At 99 minutes: utility room dryer 1 stopped.
-At 100 minutes: utility room dryer 1 running with dryness level 1.
-QT4-3 Temporal-Conflict Start dryer 1 in the bathroom at twelve thirty six PM.
-Pause dryer 1 in the bathroom immediately when dryer 1 in the utility room finishes to avoid tripping the breaker and keep the laundry loads in order.
-None At 43 minutes: bathroom dryer 1 running with dryness level 1; at 44 minutes: bathroom dryer 1 paused.
-A.6 LLM Judge Validation
-To validate the LLM-based judging, we compared its assessments to human labels on a random subset of 70 episodes spanning all judge-scored tasks.
-Human annotators showed very high inter-rater reliability (Cohen’s $\kappa$ = 0.913).
-The LLM-Judge achieved substantial agreement with the consensus human labels (Cohen’s $\kappa$ = 0.826).
-These results support using the LLM-Judge as a reliable substitute for human evaluation in our benchmark.
-After manually reviewing the 155 cases that the LLM-Judge evaluated as incorrect, we found that only 5 were misclassifications, underscoring the reliability of the evaluation.
-The detailed error distributions can be found in Table  A.11.
-A.7 Experimental Setup
-We evaluate 11 recent closed and open-source models: Llama-4-Scout (Meta AI, 2024), Llama-4-Maverick (Meta AI, 2024), Qwen3-32B (Yang et al., 2025), Qwen3-235B-A22B (Yang et al., 2025), Gemma-3-12B-it (Team et al., 2025), Gemma-3-27B-it (Team et al., 2025), Gemini-2.5-Flash (Comanici et al., 2025), Gemini-2.5-Flash-Lite (Comanici et al., 2025), GPT-4.1, GPT-4.1-mini (OpenAI, 2025a), and GPT-4.1-nano (OpenAI, 2025a).
-All models were accessed via the OpenRouter API (OpenRouter, 2025) to ensure standardized access and comparability.
-A.8 ReAct Prompt
-ReAct Prompt
-A.9 LLM Judge Prompt
-This section presents examples of the LLM Judge Prompts that were used to evaluate smart home LLM agents. Each query type (QT) has a dedicated evaluation prompt with specific criteria.
-A.9.1 QT1 Feasible Judge Prompt
-QT1 Feasible Judge Prompt (Normal)
-A.9.2 QT1 Infeasible Judge Prompt
-QT1 Infeasible Judge Prompt
-A.9.3 QT2 Infeasible Judge Prompt
-QT2 Infeasible Judge Prompt
-A.9.4 QT2 Infeasible-Nonexistence Judge Prompt
-QT2 Infeasible Judge Prompt
-A.9.5 QT3 Infeasible Judge Prompt
-QT3 Infeasible Judge Prompt
-A.9.6 QT4-1 Judge Prompt
-QT4-1 Judge Prompt
-A.9.7 QT4-2 Judge Prompt
-QT4-2 Judge Prompt
-A.9.8 QT4-3 Judge Prompt
-QT4-3 Judge Prompt
-A.10 Error Taxonomy Details
-Table 7: Error Types in Feasible Episodes
-Error Type Definition Example
-Environment Perception Errors (EP) Failure to correctly perceive or retrieve a value of environmental variables. Querying wrong sensor, misidentifying device state, guessing instead of perceiving.
-Intent Inference Errors (II) Misinterpreting user’s underlying goal. Not executing actual commands even when a user’s intention is clear.
-Device Control Errors (DC) Executing the wrong device, wrong command, or missing control steps. setting wrong channel, adjusting fan speed without turning it on first.
-Action Planning Errors (AP) Incorrect or incomplete construction of the control workflow. Breaking logical dependencies, only executing part of a multi-goal query without consideration.
-Temporal Reasoning Errors (TR) Miscalculating relative/absolute times or sequence alignment. Scheduling “in 10 minutes” at wrong time, miscomputing dishwasher completion.
-Table 8: Error Types in Infeasible Episodes
-Error Type Definition Example
-Contradiction Mishandling Errors (CM) The agent detects a contradiction but does not follow the proper instruction-following rule.
-e.g., instead of informing the user that a requested action is impossible, it arbitrarily manipulates other devices or ignores the instruction.
-Contradiction Blindness Errors (CB) The agent completely fails to recognize a contradiction and executes the request as if it were valid.
-e.g., dimming an on/off light, scheduling conflicting temporal actions without noticing inconsistency.
-LLM-Judge Errors (LJ) Errors caused not by the agent but by the evaluation system misclassifying or overlooking behavior.
-e.g., penalizing an informative refusal as a failure, or wrongly accepting hallucinated control as valid.
-A.11 Error Type Distributions
-Error Type QT2 QT3 QT4-1 QT4-2 QT4-3
-Environment Perception (EP) 3 0 4 1 0
-Intent Inference (II) 3 1 0 4 5
-Device Control (DC) 20 7 13 13 8
-Action Planning (AP) 2 0 6 3 7
-Temporal Reasoning (TR) 0 0 2 6 13
-Total 28 8 25 27 33
-Table 9: Error type distribution of GPT-4.1 in feasible episodes.
-Error Types QT1 QT2 QT3 QT4-1 QT4-2 QT4-3
-Contradiction Mishandling (CM) 8 24 6 5 6 1
-Contradiction Blindness (CB) 0 5 0 40 25 30
-LLM-Judge (LJ) 1 0 0 0 1 2
-Total 9 29 6 45 32 33
-Table 10: Error type distribution of GPT-4.1 in infeasible episodes.
-A.12 Equations of Aggregators
-$S_{r,t+1}=S_{r,t}+\sum_{d\in D_{S,r}}\Delta S_{d,r}(t),$ (1)
-where $D_{S,r}$ denotes the set of devices in room $r$ that are defined to affect state $S$, and $\Delta S_{d,r}(t)$ represents the contribution of device $d$ at tick $t$ to $S$ in room $r$.
-A.13 QT3 Error Type Distribution
-Figure 6: Distribution of tool execution error types in successful QT3 Feasible cases across four models.
-A.14 LLM Usage
-We disclose that LLMs (OpenAI GPT-5) were used as a general-purpose assistant for language editing, including polishing wording, improving clarity, and maintaining consistency.
-LLMs were not involved in research design, implementation, experiments, or analysis.
-All scientific contributions and claims in this paper are the sole responsibility of the authors.
+[…]
+Name Description Args
+[…]